# Finite Neural Networks as Mixtures of Gaussian Processes: From Provable Error Bounds to Prior Selection

### v1

@@ 7.2 Construction of a SNN Approximation of a GMM
    $\hat{{\mathbb{W}}}_{2,{R},k}$ is taken as the ${\mathbb{MW}}_{2}$ distance between $q_{nn({\mathcal{X}}),k}$ and ${R}_{M}\#q_{nn({\mathcal{X}}),k}$ (line 3 in Algorithm 1). Given the solution of the discrete optimal transport problem in Eqn. (20), the ${\mathbb{MW}}_{2}$ distance between two Gaussian mixtures reduces to the sum of 2-Wasserstein distances between Gaussian components, which have closed-form expressions (Givens and Shortt, 1984) that are differentiable with respect to the means and covariances of the components. In point 2, we showed that the means and covariances of $q_{nn({\mathcal{X}}),k}$ and ${R}_{M}\#q_{nn({\mathcal{X}}),k}$ are piecewise differentiable w.r.t the mean and covariance of $p_{{w}}$; therefore, $\hat{{\mathbb{W}}}_{2,{R},k}$ is piecewise differentiable with respect to $p_{{w}}$.
From the piecewise differentiability of $\hat{\mathcal{S}}_{2,{{w}},k}$, $\hat{{\mathbb{W}}}_{2,{\Delta},k}$, and $\hat{{\mathbb{W}}}_{2,{R},k}$ with respect to the mean and covariance of $p_{{w}}$ for all layers $k$, it naturally follows that $\hat{{\mathbb{W}}}_{2}(p^{\psi}_{nn({\mathcal{X}})},q^{\psi}_{nn({\mathcal{X}})})$ is piecewise differentiable with respect to the mean and covariance of $p_{{w}}$ using simple chain rules. For the second term of the objective, ${\mathbb{MW}}_{2}(q^{\psi}_{nn({\mathcal{X}})},q_{gp({\mathcal{X}})})$, we can apply the same reasoning from point 3 to conclude that it is piecewise differentiable with respect to the means and covariances of the components of the GMM $q^{\psi}_{nn({\mathcal{X}})}$, which, according to point 2, are piecewise differentiable with respect to the mean and covariance of $p_{{w}}$. Therefore, we can conclude that the objective in Eqn. (29) is piecewise differentiable with respect to the mean and covariances of the weight distribution of the SNN. Note that, in practice, the gradients can be computed using automatic differentiation, as demonstrated in Subsection 8.3 where we encode informative priors for SNNs via Gaussian processes by solving the optimization problem in Eqn. (29).
@@
While straightforward automatic differentiation shows to handle the discontinuity in the gradients well in practice, as analyzed in Subsection 8.3, more advanced techniques for non-smooth optimization can be employed to solve Eqn. (29) (Mäkelä, 2002; Burke et al., 2020).
@@ 8 Experimental Results
We consider networks composed of fully connected and convolutional layers, among which the VGG style architecture (Simonyan and Zisserman, 2014). We denote by $[n_{l}\times n_{n}]$ an architecture with $n_{l}$ fully connected layers, each with $n_{n}$ neurons. The composition of different layer types is denoted using the summation sign. For each SNN, we report the inference techniques used to train its stochastic layers in brackets, while deterministic layers are noted without brackets. For instance, VGG-2 + [1x128] (VI) represents the VGG-2 network with deterministic weights, stacked with a fully connected stochastic network with one hidden layer of 128 neurons learned using Variational Inference (VI). For regression tasks, the networks use tanh activation functions, whereas ReLU activations are employed for classification tasks. For VI inference, we use VOGN (Khan et al., 2018), and for deterministic and dropout networks, we use Adam.
@@
$\overline{W}_{2}(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})})\coloneqq\frac{{\mathbb{W}}_{2}(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})})}{{\mathbb{E}}_{{{z}}\sim q_{nn({\mathcal{X}})}}[\|{{z}}\|^{2}]^{\frac{1}{2}}},$
@@
which we refer to as the relative 2-Wasserstein distance. We use $\overline{W}_{2}$ because, as guaranteed by Lemma 2, it provides an upper bound on the relative difference between the second moment of $p_{nn({\mathcal{X}})}$ and $q_{nn({\mathcal{X}})}$, i.e.,
$\frac{\big{|}{\mathbb{E}}_{{{z}}\sim p_{nn({\mathcal{X}})}}[\|{{z}}\|^{2}]^{\frac{1}{2}}-{\mathbb{E}}_{\tilde{{{z}}}\sim q_{nn({\mathcal{X}})}}[\|\tilde{{{z}}}\|^{2}]^{\frac{1}{2}}\big{|}}{{\mathbb{E}}_{\tilde{{{z}}}\sim q_{nn({\mathcal{X}})}}[\|\tilde{{{z}}}\|^{2}]^{\frac{1}{2}}}\leq\overline{W}_{2}(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})}).$
@@
Note that in our framework, $q_{nn({\mathcal{X}})}$ is a GMM, hence ${\mathbb{E}}_{{{z}}\sim q_{nn({\mathcal{X}})}}[\|{{z}}\|^{2}]$ can be computed analytically.141414 We have that ${\mathbb{E}}_{{{z}}\sim q_{nn({\mathcal{X}})}}[\|{{z}}\|^{2}]=\|{m}\|^{2}+\mathrm{trace}(\Sigma)$, where ${m}$ and $\Sigma$ are the mean vector and covariance matrix of $q_{nn({\mathcal{X}})}$, respectively. Closed-forms for ${m}$ and $\Sigma$ exist for $q_{nn({\mathcal{X}})}\in\mathcal{G}_{<\infty}({\mathbb{R}}^{n})$. In our experiments, we report both a formal bound of $\overline{W}_{2}(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})})$ obtained using Theorem 13 and an empirical approximation. To compute the empirical approximation of $\overline{W}_{2}(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})})$, we use $31e$ 3 Monte Carlo (MC) samples from $p_{nn({\mathcal{X}})}$ and $q_{nn({\mathcal{X}})}$ and solve the resulting discrete optimal transport problem to approximate ${\mathbb{W}}_{2}(p_{nn({{x}})},q_{nn({{x}})})$.
Dataset Network Architecture Emp. Formal
NoisySines [1x64] (VI) 0.00109 0.33105
[1x128] (VI) 0.00130 0.13112
[2x64] (VI) 0.00355 1.63519
[2x128] (VI) 0.01135 2.48488
Kin8nm [1x128] (VI) 0.00012 0.04243
[2x128] (VI) 0.00018 0.46303
MNIST [1x128] (VI) 0.00058 0.05263
[2x128] (VI) 0.00127 0.44873
Fashion MNIST [1x128] (VI) 0.00017 0.06984
[2x128] (VI) 0.02787 1.45466
CIFAR-10 VGG-2 + [1x128] (VI) 0.00617 0.57670
VGG-2 + [2x128] (VI) 0.04901 2.35686
MNIST [1x64] (Drop) 0.27501 0.94371
[2x64] (Drop) 0.41874 1.37594
CIFAR-10 VGG-2 (Drop) + [1x32] (VI) 0.27492 179.9631
@@
Table 2: The empirical estimates and formal bounds on the relative 2-Wasserstein distance between various SNNs and their Gaussian Mixture approximates obtained via Algorithm 1 using signature approximations of size $10$ and a compression size (M) of $5$. For the dropout networks, instead of Algorithm 3, we use the compression procedure as described in Appendix B. The reported values are the average over $100$ randomly selected points from the test sets.
@@ 8.1 Gaussian Mixture Approximations of SNNs
In this subsection, we experimentally evaluate the effectiveness of Algorithm 1 in constructing a GMM that is $\epsilon$-close to the output distribution of a given SNN. We first demonstrate that, perhaps surprisingly, even a small number of discretization points for the signature operation and a large reduction in the compression operation (i.e., $M$ relatively small), often suffice for Algorithm 1 to generate GMMs that accurately approximate SNNs both empirically and formally. Then, we provide empirical evidence that by increasing the signature and compression sizes, the approximation error of the GMM can be made arbitrarily small and the formal error upper bound resulting from Theorem 13 converges to 00.
@@
We start our analysis with Table 2, where we assess the performance of Algorithm 1 in generating GMM approximations of SNN trained using both VI and dropout. The results show that for the SNNs trained with VI, even with only 10 signature points, Algorithm 1 is able to generate GMMs, whose empirical distance from the true BNN is always smaller than $10^{-2}$. This can be partly explained as for VI generally only a few neurons are active (Louizos et al., 2017; Frankle and Carbin, 2018), so it is necessary to focus mostly on these on the signature approximation to obtain accurate GMM approximations at the first $K$-th layers. Moreover, the propagation of the GMM approximation through the last linear layer is performed in closed form (i.e., by Eqn. (13c) for ${\mathcal{X}}=\{{{x}}\}$) and this mitigates the influence of the approximation error of the previous layers on the final approximation error. In contrast, for the dropout networks, the empirical approximation error is, on average, one order of magnitude larger. This can be explained because the input-output distribution of dropout networks is inherently highly multimodal. Thus, as we will show in the next paragraph, to obtain tighter GMM approximations, it is necessary to use $M>5$ (the value used in Table 2) and allow for a larger size in the GMM approximation.
Note that the formal error bounds tend to become more conservative with increasing network depth. This is related to the transformations in the (stochastic) linear layers, for which the (expected) norm of the weight matrices is used to bound its impact on the Wasserstein guarantees (see the linear operator bound in Table 11). It is also interesting to note how the error bounds tend to be more accurate, and consequently closer to the empirical estimates, for the dropout networks. This is because for VI it is necessary to bound the expected matrix norm of each linear operation, which has to be upper bounded using Lemma 14, thus introducing conservatism. Fortunately, as we will illustrate in the following paragraph and as we have mathematically proven in the previous sections, in all cases the error bounds can be made arbitrarily small by increasing signature size and $M$.
Figure 5: The formal bounds on the relative 2-Wasserstein distance between various SNNs and their Gaussian Mixture approximations obtained via Algorithm 1 for a compression size ($M$) of $3$ and different signature sizes. The plotted lines show the average of 100 random test points.
Uniform Convergence
We continue our analysis with Figure 5, where we conduct an extensive analysis of how the formal error bound from Theorem 13 changes with increasing the signature size for various SNN architectures trained on both regression and classification datasets. The plot confirms that with increasing signature sizes, the approximation error, measured as $\overline{W}_{2}$, tends to decrease uniformly. In particular, we observe an inverse linear rate of convergence, which is consistent with the theoretical asymptotic results established in Graf and Luschgy (2007) on the convergence rates of the Wasserstein distance from signatures for increasing signature size. [^15] In Appendix C, we experimentally investigate in more detail the uniform convergence of the 2-Wasserstein distance resulting from the signature operation for Gaussian mixtures of different sizes.
@@
Table 3: The formal bounds on the relative 2-Wasserstein distance between various SNNs and their Gaussian Mixture approximates obtained via Algorithm 1 for a signature size of $100$ and different compression sizes (M). The reported values are the average of 100 random test points.
@@
Figure 6: Kernels and means for various SNNs obtained via Monte Carlo sampling (left two columns) and of our GMM approximation (right two columns) for $30$ input points on MNIST in (I) and CIFAR-10 in (II) and (III). Instead of the full kernel matrices, which have dimensionality $300\times 300$, we store the trace of the sub-blocks in the output dimension, and show the $30\times 30$ matrix that captures the covariance between the input points. The rows and columns are grouped according to the classes, where the colored regions on the y-axis in the mean plots mark the classes. (I) shows the approximate kernels and the predictive means for a fully connected SNN trained via VI, which gives 98% test accuracy. We see in the kernel that examples with the same class labels are correlated. In (II), the VGG network with dropout is combined with a linear stochastic layer on the more complex CIFAR-10 data set, where we obtain 65% accuracy. (III) shows VGG stacked with a linear layer to which we both apply dropout and obtain 64% accuracy on CIFAR-10.

### v2

@@ 7.2 Construction of a SNN Approximation of a GMM
    ${\mathbb{MW}}_{2}\left(q_{nn({{x}}),k},{G}_{{M_{k}}}(q_{nn({{x}}),k})\right)$ is taken as the ${\mathbb{MW}}_{2}$ distance between $q_{nn({\mathcal{X}}),k}$ and ${G}_{{M_{k}}}(q_{nn({\mathcal{X}}),k})$ (line 3 in Algorithm 1). Given the solution of the discrete optimal transport problem in Eqn (2), the ${\mathbb{MW}}_{2}$ distance between two Gaussian mixtures reduces to the sum of 2-Wasserstein distances between Gaussian components, which have closed-form expressions (Givens and Shortt, 1984) that are differentiable with respect to the means and covariances of the components. In point 2, we showed that the means and covariances of $q_{nn({\mathcal{X}}),k}$ and ${G}_{{M_{k}}}(q_{nn({\mathcal{X}}),k})$ are piecewise differentiable w.r.t the mean and covariance of $p_{{w}}$; therefore, ${\mathbb{MW}}_{2}\left(q_{nn({{x}}),k},{G}_{{M_{k}}}(q_{nn({{x}}),k})\right)$ is piecewise differentiable with respect to $p_{{w}}$.
From the piecewise differentiability of all three quantities with respect to the mean and covariance of $p_{{w}}$ for all layers $k$, it naturally follows that $\hat{{\mathbb{W}}}_{2}\big{(}p^{\psi}_{nn({\mathcal{X}})},q^{\psi}_{nn({\mathcal{X}})}\big{)}$ is piecewise differentiable with respect to the mean and covariance of $p_{{w}}$ using simple chain rules. For the second term of the objective, ${\mathbb{MW}}_{2}\big{(}q^{\psi}_{nn({\mathcal{X}})},q_{gp({\mathcal{X}})}\big{)}$, we can apply the same reasoning from point 3 to conclude that it is piecewise differentiable with respect to the means and covariances of the components of the GMM $q^{\psi}_{nn({\mathcal{X}})}$, which, according to point 2, are piecewise differentiable with respect to the mean and covariance of $p_{{w}}$. Therefore, we can conclude that the objective in Eqn (33) is piecewise differentiable with respect to the mean and covariances of the parameter distribution of the SNN. Note that, in practice, the gradients can be computed using automatic differentiation, as demonstrated in Subsection 8.3 where we encode informative priors for SNNs via Gaussian processes by solving the optimization problem in Eqn (33).
@@
While straightforward automatic differentiation shows to handle the discontinuity in the gradients well in practice, as analyzed in Subsection 8.3, more advanced techniques for non-smooth optimization can be employed to solve Eqn (33) (Mäkelä, 2002; Burke et al., 2020).
@@ 8 Experimental Results
We consider networks composed of fully-connected and convolutional layers, among which the VGG style architecture (Simonyan and Zisserman, 2014). We denote by $n_{l}\times n_{n}]$ an architecture with $n_{l}$ fully-connected layers, each with $n_{n}$ neurons. The composition of different layer types is denoted using the summation sign. For each SNN, we report the inference techniques used to train its stochastic layers in brackets, while deterministic layers are noted without brackets. For instance, VGG-2 + [1x128] (VI) represents the VGG-2 network with deterministic parameters, stacked with a fully-connected stochastic network with one hidden layer of 128 neurons learned using Variational Inference (VI). For regression tasks, the networks use tanh activation functions, whereas ReLU activations are employed for classification tasks. For VI inference, we use VOGN (Khan et al., 2018), and for deterministic and dropout networks, we use Adam.
@@
$\overline{W}_{2}\left(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})}\right)\coloneqq\frac{{\mathbb{W}}_{2}\left(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})}\right)}{{\mathbb{E}}_{{{z}}\sim q_{nn({\mathcal{X}})}}[\|{{z}}\|^{2}]^{1/2}},$ (34)
@@
which we refer to as the relative 2-Wasserstein distance. We use $\overline{W}_{2}$ because, as guaranteed by Lemma 28 in the appendix, it provides an upper bound on the relative difference between the second moment of $p_{nn({\mathcal{X}})}$ and $q_{nn({\mathcal{X}})}$, i.e.,
$\frac{\big{|}{\mathbb{E}}_{{{z}}\sim p_{nn({\mathcal{X}})}}[\|{{z}}\|^{2}]^{1/2}-{\mathbb{E}}_{\tilde{{{z}}}\sim q_{nn({\mathcal{X}})}}[\|\tilde{{{z}}}\|^{2}]^{1/2}\big{|}}{{\mathbb{E}}_{\tilde{{{z}}}\sim q_{nn({\mathcal{X}})}}[\|\tilde{{{z}}}\|^{2}]^{1/2}}\leq\overline{W}_{2}\left(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})}\right).$ (35)
@@
Note that in our framework, $q_{nn({\mathcal{X}})}$ is a GMM, hence ${\mathbb{E}}_{{{z}}\sim q_{nn({\mathcal{X}})}}[\|{{z}}\|^{2}]$ can be computed analytically.171717 We have that ${\mathbb{E}}_{{{z}}\sim q_{nn({\mathcal{X}})}}[\|{{z}}\|^{2}]=\|{m}\|^{2}+\mathrm{trace}(\Sigma)$, where m{m} and $\Sigma$ are the mean vector and covariance matrix of $q_{nn({\mathcal{X}})}$, respectively. Closed-forms for m{m} and $\Sigma$ exist for $q_{nn({\mathcal{X}})}\in\mathcal{G}_{<\infty}({\mathbb{R}}^{n})$. In our experiments, we report both a formal bound of $\overline{W}_{2}\left(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})}\right)$ obtained using Theorem 17 and an empirical approximation. To compute the empirical approximation of $\overline{W}_{2}\left(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})}\right)$, we use $1e3$ Monte Carlo (MC) samples from $p_{nn({\mathcal{X}})}$ and $q_{nn({\mathcal{X}})}$ and solve the resulting discrete optimal transport problem to approximate ${\mathbb{W}}_{2}\left(p_{nn({{x}})},q_{nn({{x}})}\right)$.
Grid Cross
Dataset Network Architecture Emp. Formal Emp.
NoisySines [1x64] (VI) 0.00109 0.33105 0.00028
[1x128] (VI) 0.00130 0.13112 0.00041
[2x64] (VI) 0.00355 1.63519 0.00305
[2x128] (VI) 0.01135 2.48488 0.00956
Kin8nm [1x128] (VI) 0.00012 0.04243 0.00012
[2x128] (VI) 0.00018 0.46303 0.00005
MNIST [1x128] (VI) 0.00058 0.05263 0.00037
[2x128] (VI) 0.00127 0.44873 0.00124
Fashion MNIST [1x128] (VI) 0.00017 0.06984 0.00013
[2x128] (VI) 0.02787 1.45466 0.02228
CIFAR-10 VGG-2 + [1x128] (VI) 0.00617 0.57670 0.00363
VGG-2 + [2x128] (VI) 0.04901 2.35686 0.02890
MNIST [1x64] (Drop) 0.27501 0.94371 0.18400
[2x64] (Drop) 0.41874 1.37594 0.08603
CIFAR-10 VGG-2 (Drop) + [1x32] (VI) 0.27492 179.9631 0.18996
@@
Table 2: The empirical estimates and formal bounds on the relative 2-Wasserstein distance between various SNNs and their Gaussian Mixture approximates obtained via Algorithm 1 using signature approximations of size $10$ and a compression size ($M_{k}$) of $5$ for all layers. The Grid column reports the error for approximates constructed using Algorithm 2. The Cross column reports the empirical errors for approximates constructed using the alternative cross-shaped signatures via Algorithm 4, for which formal bounds are intractable for this method. For the dropout networks, instead of Algorithm 3, we use the compression procedure as described in Appendix E. The reported values are the average over $100$ randomly selected points from the test sets.
@@ 8.1 Gaussian Mixture Approximations of SNNs
In this subsection, we experimentally evaluate the effectiveness of Algorithm 1 in constructing a GMM that is $\epsilon$-close to the output distribution of a given SNN. We first demonstrate that, perhaps surprisingly, even a small number of discretization points for the signature operation and a large reduction in the compression operation (i.e., $M_{k}$ relatively small), often suffice for Algorithm 1 to generate GMMs that accurately approximate SNNs both empirically and formally. Then, we provide empirical evidence that by increasing the signature and compression sizes, the approximation error of the GMM can be made arbitrarily small and the formal error upper bound resulting from Theorem 17 converges to 0.
@@
We start our analysis with Table 2, where we assess the performance of Algorithm 1 in generating GMM approximations of SNN trained using both VI and dropout. The results show that for the SNNs trained with VI, even with only 10 signature points per-component, Algorithm 1 is able to generate GMMs, whose empirical distance from the true BNN is always smaller than $10^{-2}$. This can be partly explained as for VI generally only a few neurons are active (Louizos et al., 2017; Frankle and Carbin, 2018), so it is necessary to focus mostly on these on the signature approximation to obtain accurate GMM approximations at the first $K$-th layers (See Remark 22). Moreover, the propagation of the GMM approximation through the last fully-connected layer is performed in closed form (i.e., by Eqn (19c) for ${\mathcal{X}}=\{{{x}}\}$) and this mitigates the influence of the approximation error of the previous layers on the final approximation error. In contrast, for the dropout networks, the empirical approximation error is, on average, one order of magnitude larger. This can be explained because the input-output distribution of dropout networks is inherently highly multimodal. Thus, as we will show in the next paragraph, to obtain tighter GMM approximations, it is necessary to use $M>5$ (the value used in Table 2) and allow for a larger size in the GMM approximation.
@@
Table 3: The formal bounds on the relative 2-Wasserstein distance between various SNNs and their Gaussian Mixture approximates obtained via Algorithm 1 for a signature size of $100$ and different compression sizes ($M_{k}$). The reported values are the average of 100 random test points.
Note that the formal error bounds tend to become more conservative with increasing network depth. This is related to the transformations in the (stochastic) fully-connected layers, for which the (expected) norm of the weight matrices is used to bound its impact on the Wasserstein guarantees (see the affine operator bound in Table 13). It is also interesting to note how the error bounds tend to be more accurate, and consequently closer to the empirical estimates, for the dropout networks. This is because for VI it is necessary to bound the expected matrix norm of each affine operation, which has to be upper bounded (see Appendix C), thereby introducing conservatism. Fortunately, as we will illustrate in the following paragraph and as we have mathematically proven in the previous sections, in all cases the error bounds can be made arbitrarily small by increasing signature size and $M$.
Lastly, in settings where formal error bounds are not required, alternative signature-placement methods such as Algorithm  4 in the appendix, which places the signature locations in a cross along the principal axes of the components’ covariance matrix, provide a practical alternative. In the cases considered, this cross-shaped placement consistently attains higher empirical accuracy with the same number of signature points and lower runtime, so fewer points often suffice, however, it does not yield formal guarantees. The compute cost of Algorithm 1 is modest, with processing a single test point typically taking on the order of tenths of a second. [^18] All experiments were run on an Intel Core i7-10610U (1.80–2.30 GHz) with 16 GB RAM.
Uniform Convergence
We continue our analysis with Figure 6, where we conduct an extensive analysis of how the formal error bound from Theorem17 changes with increasing signature size for various SNN architectures trained on both regression and classification datasets. The plot confirms that, with increasing signature sizes, the approximation error—measured as $\overline{W}_{2}$—tends to decrease uniformly. In particular, we observe an inverse‑linear convergence rate. This behaviour matches our theoretical result in Theorem 21. That theorem implies that a per‑dimension grid size of order $N=O(\epsilon^{-1}\sqrt{\log(1/\epsilon)})$ is sufficient to reach a target precision $\epsilon$. [^19] In Appendix F, we experimentally investigate in more detail the uniform convergence of the 2‑Wasserstein distance resulting from the signature operation for Gaussian mixtures of different sizes.
@@
Figure 6: The formal bounds on the relative 2-Wasserstein distance between various SNNs and their Gaussian Mixture approximations obtained via Algorithm 1 for a compression size ($M_{k}$) of $3$ and different signature sizes. The plotted lines show the average of 100 random test points.
@@
Figure 7: Kernels and means for various SNNs obtained via Monte Carlo sampling (left two columns) and of our GMM approximation (right two columns) for $30$ input points on MNIST in (I) and CIFAR-10 in (II) and (III). Instead of the full kernel matrices, which have dimensionality $300\times 300$, we store the trace of the sub-blocks in the output dimension, and show the $30\times 30$ matrix that captures the covariance between the input points. The rows and columns are grouped according to the classes, where the colored regions on the y-axis in the mean plots mark the classes. (I) shows the approximate kernels and the predictive means for a fully-connected SNN trained via VI, which gives 98% test accuracy. We see in the kernel that examples with the same class labels are correlated. In (II), the VGG network with dropout is combined with a linear stochastic layer on the more complex CIFAR-10 data set, where we obtain 65% accuracy. (III) shows VGG stacked with a linear layer to which we both apply dropout and obtain 64% accuracy on CIFAR-10.